a 500 cm rope under a tension of 70 N is set into oscillation. the mass density of the rope is 160 g/cm. what is the frequency of the first harmonic mode (first overtone)?

0.42 Hz
0.132 Hz
0.0042 Hz
2.19 Hz
Answer #1

Frequency of first harmonic mode,

F2 = (2/2l)sqrt(T/u)

F2 = (1/5) sqrt(70/16) = 0.42 Hz

Comment in case any doubt please rate my answer....